OpenStudy (anonymous):

I don't understand this problem, can someone help me to please? For the quadratic equation 8x2 - 55x = -42 give the values of A,B, and C.

OpenStudy (anonymous):

Ok. A quadratic equation is in the form \(ax^{2} + bx + c = 0\).Therefore, what are a, b, and c in this quadratic equation? \(8x^{2} - 55x + 42 = 0\)?

OpenStudy (anonymous):

Calcmathlete is right,,

OpenStudy (anonymous):

@Calcmathlete So is this correct ? A = 8, B = -5, C = 42

OpenStudy (anonymous):

@cornitodisc So is this correct ? A = 8, B = -5, C = 42

OpenStudy (anonymous):

Not quite. A = 8 B = -55 and C = 42

OpenStudy (anonymous):

Maybe a typo?

OpenStudy (anonymous):

yes indeed,,but B= -55

OpenStudy (anonymous):

here' an example for your question Quadratic Function A quadratic function of the variable x is a function that can be written in the form f(x) = ax2 + bx + c Function form y = ax2 + bx + c Equation form where a, b, and c are fixed numbers (with a ≠ 0). Examples 1. f(x) = 3x2 - 2x + 1 a = 3, b = -2, c = 1 2. g(x) = -x2 a = -1, b = 0, c = 0 3. h(x) = 3x + 1 Not a quadratic function because a = 0